Question: "Determine the genotypic frequencies of offspring for a genetic trait with two alleles, R and r, with the previous generation's allelic frequencies of R=0.1 and r=0.9."

My professor said that you can use the formula p^2+2pq+q^2=1 where R=p and r=q, but i'm not entirely sure how to go about this.

-Population practicies random mating (no picking your mates)

-No immigrration or emigration (i.e., a closed population)
Answer #1

Allele Frequencies Equation:-

(p + q)2 =p2 + 2pq + q2

It must also be noted that the sum of all the allelic frequencies and of all the genotypic frequencies must always be 1. The Hardy-Weinberg law is applicable when there is random mating.

(0.1)^2 + 2(0.1*0.9) + (0.9)^2=1

0.01 + 2(0.9) + 0.81=1

0.01 + 1.8 + 0.81=1

2.62=1

0.62